OpenStudy (anonymous):

which follow these converge? Also calculate 1. F0 = F1 = 1, and Fn + 1 = Fn + Fn-1 2. a1 = 2 and a n + 1 = (to) ^ (1/2) 3. A0 = 1 and an + 1 = an + m ^ n for every m∈R

OpenStudy (xapproachesinfinity):

the first is Fibonacci in recurrence formula

OpenStudy (xapproachesinfinity):

haha go sleep at any rate how do you decide convergence?

OpenStudy (xapproachesinfinity):

yes diverge since the ratio goes to psi which greater than 1

OpenStudy (xapproachesinfinity):

psi=golden mean

OpenStudy (xapproachesinfinity):

hmm ya you can say increasing without bound

OpenStudy (xapproachesinfinity):

limit is another way to go at it, that's i asked what does the asker know about convergence

OpenStudy (xapproachesinfinity):

haha ya, i never memorized

OpenStudy (xapproachesinfinity):

hmm well good that you mentioned boundness

OpenStudy (xapproachesinfinity):

okay!

OpenStudy (xapproachesinfinity):

hmm lets the same way loser did the first a1=2 a2=root2 a3=root (root2) so is this bounded or not

OpenStudy (xapproachesinfinity):

what language is that problem was written?

OpenStudy (anonymous):

Deutsch

OpenStudy (xapproachesinfinity):

hmm i see. what level of mathematics are you doing?

OpenStudy (xapproachesinfinity):

see that our terms are getting closer and closer to zero yes 2>root2>root(root(2))>...........

OpenStudy (xapproachesinfinity):

so a_n+1<=a_n and bounded therefore converges

OpenStudy (anonymous):

sorry , i do not speaking well english .. I am in my first year at the college

OpenStudy (xapproachesinfinity):

oh i see, no problem but you understand the math yes

OpenStudy (xapproachesinfinity):

\[a_1=2\\a_2=\sqrt{2}\\a_3=\sqrt{\sqrt{2}}\\.\\.\\.\\a_{n+1}=\sqrt{a_{n}}\] note that \[a_{n+1}\leq a_n\] and \[a_n \] is bounded so it must converges

OpenStudy (xapproachesinfinity):

hmm there is little problem @perl the sequence is not non increasing yes?

OpenStudy (perl):

not non increasing means decreasing?

OpenStudy (xapproachesinfinity):

not really isn't non increasing assume that terms might equal

OpenStudy (perl):

we have a sequence $$ \LARGE 2, 2^{\frac 12}, 2^{\frac14},2^{\frac 18},.. 2^{\frac 1{2^{n}}},... $$

OpenStudy (xapproachesinfinity):

that's my question this sequence is strictly decreasing

OpenStudy (perl):

right

OpenStudy (perl):

it is bounded above by 2, bounded below by 0 (or 1), and it is monotonic therefore there exists a greatest lower bound

OpenStudy (perl):

and that is the limit of the sequence

OpenStudy (perl):

it is bounded below by zero since all the terms in the sequence are positive

OpenStudy (perl):

and bounded above by 2 since it is decreasing, and 2 is the first term

OpenStudy (perl):

we have a sequence $$ \LARGE 2 \geq 2^{\frac 12} \geq 2^{\frac14} \geq 2^{\frac 18},.. \geq2^{\frac 1{2^{n}}},... $$

OpenStudy (xapproachesinfinity):

yeah i wanted to use boundness theorem but i thought it has to be nonincreasing

OpenStudy (perl):

In the mathematical field of real analysis, the monotone convergence theorem is any of a number of related theorems proving the convergence of monotonic sequences (sequences that are increasing or decreasing) that are also bounded. Informally, the theorems state that if a sequence is increasing and bounded above by a supremum, then the sequence will converge to the supremum; in the same way, if a sequence is decreasing and is bounded below by an infimum, it will converge to the infimum http://en.wikipedia.org/wiki/Monotone_convergence_theorem

OpenStudy (xapproachesinfinity):

\(\color{blue}{\text{Originally Posted by}}\) @xapproachesinfinity see that our terms are getting closer and closer to zero yes 2>root2>root(root(2))>........... \(\color{blue}{\text{End of Quote}}\) that's what i was doing but had doubt

OpenStudy (perl):

the monotonic sequences convergence theorem does not have to be strictly decreasing or increasing. monotonicity is enough

OpenStudy (xapproachesinfinity):

oh ok

OpenStudy (perl):

non decreasing (sometimes called increasing) non increasing (sometimes called decreasing) http://en.wikipedia.org/wiki/Monotonic_function

OpenStudy (perl):

love your pic by the way , i forgot who you were

OpenStudy (xapproachesinfinity):

okay i see

OpenStudy (xapproachesinfinity):

poincare is awesome haha

OpenStudy (xapproachesinfinity):

watched some videos about math history, this guy is quite something

OpenStudy (xapproachesinfinity):

check nj wilberger math history one of the videos talk about poincare and other mathematicians

OpenStudy (perl):

so in a nutshell the monotonic sequence theorem says 1. If a sequence of real numbers is strictly increasing (or not decreasing) and bounded above, then its supremum is the limit. 2. If a sequence of real numbers is strictly decreasing (or not increasing) and bounded below, then its infimum is the limit. In our case we use part 2. but we have to show that an+1 < an

OpenStudy (perl):

show that \( \LARGE 2^{\frac {1}{n+1} } < 2^\frac 1n \)

OpenStudy (xapproachesinfinity):

i see thanks :)

OpenStudy (perl):

you can show it either by induction or how about raise both sides to the power of n+1

OpenStudy (xapproachesinfinity):

hmm we show that f(x)=1/2^x is decreasing so 2^(1/n+1)<2(1/n)

OpenStudy (xapproachesinfinity):

hmm i went too far induction is good actually

OpenStudy (perl):